Information from the financial statements of Ames Fabricators, Inc., included the following:
  

December 31
2018 2017
Common shares 100,000 100,000
Convertible preferred shares
(convertible into 46,000 shares of common)
20,400 20,400
10% convertible bonds
(convertible into 20,000 shares of common)
$ 1,000,000 $ 1,000,000


Ames’s net income for the year ended December 31, 2018, is $640,000. The income tax rate is 40%. Ames paid dividends of $5 per share on its preferred stock during 2018.

Required:

Compute basic and diluted earnings per share for the year ended December 31, 2018. (Enter your answers in thousands.)

Numerator / Denominator = Earnings per Share
Basic EPS / =
Diluted EPS / =
Answer #1

ANSWER:

Basic EPS=(Net Income-Preferred Dividend)/(Average Outstanding Shares)

Income for the year 2018=$640,000

Income Tax =40%*640000=$256,000

Net Income =(640000-256000)=$384,000

Preferred dividend =$5*20400=$102000

Average outstanding Shares=100000

Basic EPS=(384000-102000)/100000

=282000/100000=$2.82

Diluted EPS==(Net Income-Preferred Dividend)/(Average Outstanding Shares+ Diluted Shares)

Diluted Shares=46000+20000=66000

Diluted EPS=282000/(100000+66000)=282000/166,000=$1.69

Numerator/

Denominator

=

Earning Per Share

Basic EPS

282000

100000

=

$2.82

Diluted EPS

282000

166,000

=

$1.69

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************